bob is seven years older than al.chuck is twice as old as al.if the sum of all their ages is 75,then find the age of each person.
Answer by Leaf W.(135)

Let us call Al's age in years "x." If Bob is seven years older than Al, then we can say that Bob's age in years is x + 7 years. Also, if Chuck is twice as old as Al, that means that the age of Chuck in years is two times the age of Al in years. As a result, we can say that Chuck's age in years is 2x. To simplify, we now have the following:
x = Al's age in years
x + 7 = Bob's age in years
2x = Chuck's age in years
The problem next states that the sum (addition) of all their ages is 75, so we can say Al's age + Bob's age + Chuck's age = 75, or x + x + 7 + 2x = 75.
Now we can add up the like terms: x + x + 2x = 4x, so 4x + 7 = 75
Subtract 7 from both sides: 4x = 68
Divide both sides by 4: x = 17
Since x means Al's age, Al is 17 years old. Bob's age, x + 7, is 17 + 7 = 24 years old, and Chuck's age, 2x, is 2 * 17 = 34 years old. Therefore,
Al's age: 17 years
Bob's age: 24 years
Chuck's age: 34 years
Hope I helped! =)